OpenStudy (anonymous):

\[\sqrt{8}-\sqrt{2}\]

OpenStudy (anonymous):

=sqrt2 = 1.414

OpenStudy (mertsj):

\[\sqrt{8}=\sqrt{4\times2}=2\sqrt{2}, and 2\sqrt{2}-1\sqrt{2}=\sqrt{2}\]